OpenStudy (anonymous):

Find the Regular Grammar for the following Regular Expression: a(a+b)*(ab*+ba*)b

geerky42 (geerky42):

@wio may know. Tagged him, so he will come here when he get on.

OpenStudy (anonymous):

With questions like these, sometimes it helps to just construct a formal grammar, and then expand it so that it is regular.

OpenStudy (anonymous):

Okay, so we have \(a(a+b)^*(ab^*+ba^*)b\). Let's suppose \(T\) represents \(a(a+b)^*(ab^*+ba^*)\), then we can start with a rule: \[ \begin{array}{lll} S &\to& Tb \end{array} \]Now we just need to find \(T\). Again, suppose \(C\) represents \(a(a+b)^*\). Consider that \(C(ab^*+ba^*)\) is equivalent to \(Cab^*+Cba^*\). First let's construct a means of doing \(Cab^*\)\[ \begin{array}{lll} N&\to& Ca \\ &|&Nb\\ \end{array} \]Next we need to consider \(Cba^*\) and join them:\[ \begin{array}{rcl} M&\to& Cb\\ &|&Ma\\ \\ T &\to& M \\ &|& N \end{array} \]Now finally for \(C\). We've already some something similar before.\[ \begin{array}{lll} C&\to& a \\ &|&Ca\\ &|&Cb \end{array} \]So now let's just list out the whole grammar:\[ \begin{array}{llll} S&\to& Tb \\ T&\to&M &|& N\\ M&\to&Cb&|& Ma\\ N&\to&Ca&|& Nb\\ C&\to&a&|&Ca &|& Cb \end{array} \]Now, I'm using what is called a left regular grammar. That means I will only use rules of the form \(\epsilon,a,Na\), where \(a\) is a terminal and \(N\) is a non-terminal that is also a left regular grammar. A right regular grammar is similar, but instead of \(Na\) you have to use rules of the form \(aN\). By restricting to these types of rules, you ensure your grammar is regular.